OpenStudy (anonymous):

Integrate (3x^2 - 28x + 50)/(x^3 - 10x^2 + 25x) dx using partial fractions method.

OpenStudy (anonymous):

\[ \frac{ A }{ x } + \frac{ B }{ x-5 } + \frac{ C }{ (x-5)^2 }\] A = 2 and C = -3/5 How do I find B?

OpenStudy (anonymous):

plug it any number of x

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

I am trying

OpenStudy (anonymous):

Can you show it to me, please?

OpenStudy (anonymous):

\[\frac{A}{x} + \frac{B}{(x-5)}+\frac{C}{(x-5)^2} = \frac{3x^2-28x+50}{x^3-10x^2+25x}\] \[\frac{A(x-5)^2+Bx(x-5)+Cx}{x(x-5)^2} = \frac{3x^2-28x+50}{x(x-5)^2}\] \[A(x^2 -10x+25) + B(x^2 - 5x) + Cx = 3x^2 - 28x + 50\] \[Ax^2 - 10Ax + 25A +Bx^2 -5Bx+Cx = 3x^2 - 28x + 50\] \[(A + B)x^2 - (10A + 5B - C)x +25A = 3x^2 - 28x + 50\] Setting the coefficients equal gives: \[A+B=3\] \[10A + 5B - C = 28\] \[25A = 50\] This gives A = 2, plugging into the first gives B = 1,so \[20 + 5 - c = 28\] C = -3 So the result is: \[\frac{2}{x} + \frac{1}{x-5} - \frac{3}{(x-5)^2}\]